U are traveling in a car at 60mph and you have driven for 8 hours with 2 stops of 15 minutes each. How far have you gone

Answer by checkley79(3341)

D=RT
D=60(8-2*1/4)
D=60(8-.5)
D=60*7.5
D=450 MILES ANS.